OpenStudy (anonymous):

Find a polynomial for the perimeter and for area. The length of the rectangle is d+3 and the width is d. The perimeter is __ The area is __

OpenStudy (anonymous):

4d + 6

OpenStudy (anonymous):

d^2 + 3d <--------area

OpenStudy (anonymous):

given l=d+3 b=d area = l*b=d(d+3)\[=d ^{2}+3d\] primeter=2(l+b) =2(d+d+3)=4d+6

OpenStudy (anonymous):

hope u have understood

OpenStudy (anonymous):

Two different answers. Now I'm confused lol

OpenStudy (anonymous):

lol they are same

OpenStudy (anonymous):

oops lol

OpenStudy (anonymous):

but both are the same answer .

OpenStudy (anonymous):

:)